OpenStudy (anonymous):

use the rational root theorem to list all possible rational roots for this equation then find any actual rational roots 2x^3-x^2+10x-5=0

OpenStudy (anonymous):

list all fractions where the numerator divides the constant 5 and the denominator divides the leading coefficient 2

OpenStudy (anonymous):

the only divisors of 5 are 5 and 1, and the divisors of 2 are 1 and 2 so the list looks like \[\pm1, \pm5,\pm\frac{1}{2},\pm\frac{5}{2}\]

OpenStudy (anonymous):

the only actual real zero is \(\frac{1}{2}\) http://www.wolframalpha.com/input/?i=2x^3-x^2%2B10x-5%3D0

OpenStudy (anonymous):

ok and is this all i do

OpenStudy (anonymous):

you are asked to find the actual zeros as well \(\frac{1}{2}\) is the only one from the list that works

OpenStudy (precal):

3rd power means that you have 3 solutions p/q helped you identify one of the solutions at 1/2 you will need to find the remaining two solutions I suggest you use synthetic division to find the other two